OpenStudy (goldrush18):

simplify 2sqrt3 +sqrt5 / sqrt3 -2sqrt5 simplify 2sqrt3 +sqrt5 / sqrt3 -2sqrt5 @Mathematics

OpenStudy (anonymous):

This doesn’t need to be simplified, it can already be solved..

OpenStudy (goldrush18):

\[2\sqrt3+\sqrt5\div \sqrt3-2\sqrt5\]

OpenStudy (goldrush18):

ok solve it please

OpenStudy (anonymous):

I believe this is what you are looking for. \[\frac{2\sqrt 3 + \sqrt 5}{\sqrt 3 - 2 \sqrt 5} \cdot \frac{\sqrt 3 + 2\sqrt 5}{\sqrt 3 + 2 \sqrt 5} = \frac{2 \cdot 3 + 2 \sqrt{15} + \sqrt{15} + 2 \cdot 5}{3-4^2\cdot 5} = \boxed{\displaystyle \frac{16+3\sqrt{15}}{-77}}\]

OpenStudy (goldrush18):

thanks im always getting confused when for example 2sqrt3*2Sqrtt5

OpenStudy (anonymous):

Whoops, that denominator should be \[3-2^2 \cdot 5 = -17\]

OpenStudy (goldrush18):

i thought that would be 4sqrt15

OpenStudy (anonymous):

Show me what you did to get that.

OpenStudy (goldrush18):

\[2\sqrt3 * 2\sqrt5=4\sqrt15\]

OpenStudy (goldrush18):

maybe im wrong i hate surds lol

OpenStudy (anonymous):

Ah yes you are definitely correct haha. Wow today must be an of day. Yes, the coefficient should be 5 instead of 3 in the numerator on the sqrt(15).

OpenStudy (goldrush18):

so the 4sqrt15 is correct?

OpenStudy (anonymous):

Yes, but it gets added to the 1sqrt15 to make 5sqrt15

OpenStudy (goldrush18):

thank you